### v1

# Seeing Sound, Hearing Sight: Uncovering Modality Bias and Conflict of AI models in Sound Localization

Abstract
Imagine hearing a dog bark and instinctively turning toward the sound—only to find a parked car, while a silent dog sits nearby. Such moments of sensory conflict challenge perception, yet humans flexibly resolve these discrepancies, prioritizing auditory cues over misleading visuals to accurately localize sounds. Despite the rapid advancement of multimodal AI models that integrate vision and sound, little is known about how these systems handle cross-modal conflicts or whether they favor one modality over another. Here, we systematicly and quantitatively investigate modality bias and conflict resolution in AI sound localization. We evaluate a wide range of state-of-the-art multimodal models and compare them against human performance in carefully designed psychophysics experiments spanning six audiovisual conditions, including congruent, conflicting, and absent visual and audio cues. Our results reveal that humans consistently outperform AI in sound localization and exhibit greater robustness to conflicting or absent visual information by effectively prioritizing auditory signals. In contrast, AI shows a pronounced bias toward vision, often failing to suppress irrelevant or conflicting visual input, leading to chance-level performance. To close this gap, we fine-tune a state-of-the-art multimodal model on our curated stereo audio-image dataset synthesized via 3D simulations. With fewer training examples, our resulting model outperforms existing AI trained on standard audio-visual benchmark datasets. Remarkably, consistent with neuroscience findings, our model exhibits a human-like horizontal localization bias—favoring left-right over up-down precision. This likely stems from training on stereo audio that mirrors the horizontal arrangement of human ears. These results highlight how the quality of sensory inputs and the physical structure of sensory systems jointly shape the precision of multimodal representation. All code, data, and models will be made public.
1 Introduction
(a) Sound source localization challenge
(b) Depth-aware stereo audio synthesis
Figure 1: (a) Sound source localization challenge in naturalistic images. The left shows a multi-source scenario with sounds from several objects (yellow boxes), while the right depicts a single-source case with only a telephone ring (red box). In both settings, the task is to localize all sounding objects based on two-channel stereo audio. To allow systematic and controllable benchmarking of human and AI performance, we focus on the single-source localization task (red), which remains challenging due to scene clutter, occlusions, and ambiguous visual cues. (b) Depth-aware stereo audio synthesis. In the 3D simulation environment, a human listener (interaural distance: 0.17m) is placed at the origin, facing the RGB image on the screen. This image and its depth image are aligned in the same direction. Using a spatial audio renderer and audio from our library, stereo sound of the target car (red box) in the RGB image can be synthesized (see Sec 2.1).
Sound source localization (SSL) is the task of identifying the spatial origin of a sound within a visual scene. It plays a fundamental role in both biological perception (van2019cortical, ) and artificial intelligence (AI) (jenni2023audio, ), enabling systems to connect what they hear with what they see. SSL requires integrating information across auditory and visual modalities to accurately determine not only what is sounding, but also where it is located. In neuroscience and psychology, SSL provides insights into how humans integrate complex multisensory information to direct attention (francl2022deep, ; oude2024triple, ) and make dynamic decisions (noppeney2021perceptual, ). In real-world practice, accurate SSL is essential for autonomous vehicles anticipating hazards (fang2022traffic, ), assistive devices guiding visually impaired users (dimas2021self, ), and interactive robots engaging in human-centered tasks (chen2025human, ; saeed2019multi, ; dworakowski2023robots, ). These applications require reliable detection of sound sources, particularly in cluttered or ambiguous scenes.
Consider a real-life example in Fig. 1(a) depicting a busy traffic intersection: cars honk, people converse, birds chirp, and an ambulance siren approaches from behind a building. Humans must rapidly detect and localize the siren despite competing sounds, partial visual occlusion, and environmental noise. Such everyday scenes highlight the challenges of SSL: resolving ambiguous visual or auditory cues, handling conflicting signals across modalities, and focusing attention on the most relevant source amidst distractions. Building AI systems capable of performing SSL tasks in these conditions remains an open problem.
Recent AI research has developed multimodal models for SSL (georgescu2023audiovisual, ; huang2023mavil, ; jia2025uni, ; li2021space, ; free, ). Methods such as contrastive learning (denseav, ; cavp, ; mo2022closer, ; qian2020multiple, ) aim to align audio-visual embeddings by maximizing the similarity between matching pairs and minimizing it between mismatched ones. Other approaches leverage cross-modal attention (mask, ; tar, ; lin2023unsupervised, ) in transformer architectures (is3, ; gao2024avsegformer, ; li2023catr, ) to allow sound features to dynamically query relevant visual regions. Some works (pv4, ; xuan2022proposal, ; senocak2022less, ) also incorporate object priors, leveraging knowledge of what typically makes sound to guide localization. However, despite these advances (gong2022contrastive, ), little is known about how such models behave when the modalities conflict or when biases emerge—such as favoring visual cues over auditory ones.
To address this gap, we systematically evaluate model behavior under six controlled audio-visual conditions: (1) Congruent — audio and visual cues align in both semantics and location; (2) conflicting visual cues, where the visual scene misleads localization; (3) absent visual cues, where the sounding object is completely occluded in the visual scene; and (4) vision-only and (5) audio-only conditions, where either vision or audio is entirely omitted. Moreover, similar to the cocktail party problem (haykin2005cocktail, ; sun2024unveiling, ), we further extend the multi-instance SSL (wang2023robust, ) to single-source stereo audio, where multiple instances of the same semantic category are present, potentially distracting the model from correctly localizing the target object. These manipulations allow us to probe how models resolve cross-modal ambiguity and characterize their reliance on each modality. To benchmark these model behaviors, we additionally conduct human behavioral studies under the same experimental conditions. Results show a clear performance gap: humans consistently outperform AI models, particularly in handling both congruent and incongruent conditions.
While numerous SSL datasets (lee2021acav100m, ; languagebind, ; ade20k, ) exist, they often suffer from limitations that hinder robust multimodal alignment in AI models. Often, these datasets (zhou2024audio, ; refavs, ; peavs, ; acappella, ; slakh, ; urbansed, ) consist of scenes with a single, large, centrally placed sounding object, making them vulnerable to visual shortcut learning, where models perform well without truly integrating audio information (cavp, ; is3, ; fnac, ). Others (vggss, ; is3, ) are constructed by pairing unrelated images and sounds from independent datasets (coco, ; imagenet, ; vggsound, ; FSDnoisy18k, ; DCASE, ; l3das22, ; vocalsound, ), leading to weak cross-modal entanglement. More recent efforts (sanguineti2021audio, ; fsd, ; aytar2016soundnet, ) synthesize audio based on physics-informed rules (mcdermott2011sound, ; gun, ) or generative AI models (zhou2020sep, ; lll, ), but they still frequently rely on mono audio, neglecting the richer spatial cues provided by stereo signals. Stereo audio, widely studied in neuroscience for its role in spatial hearing through inter-channel differences (woods2017headphone, ), remains underexplored in SSL benchmarks in AI.
To address this gap, we introduce a method that uses 3D simulation engines to generate stereo audio for images by combining standalone image and sound datasets. Unlike neuroscience approaches that require physically setting up microphones in real-world spaces, which is both time-consuming and costly, our simulation-based method offers a scalable and efficient alternative for producing spatialized audio paired with complex visual scenes. With this method, we contribute a large-scale AudioCOCO dataset, comprising 1,003,747 image-audio pairs with ground truth annotations. By simulating stereo audio that adheres to physical principles of sound propagation and spatial cues, AudioCOCO provides realistic and diverse audio-visual scenes.
To further validate the utility of our dataset, we fine-tune a state-of-the-art (SOTA) AI model on stereo audio from our AudioCOCO dataset. The resulting model outperforms existing SOTA methods, which tend to be biased toward large, centered, or vocal objects (is3, ; cavp, ; refavs, ; avs, ). Neuroscience research francl2022deep shows that humans exhibit greater precision in horizontal (left-right) sound localization than vertical (up-down). Remarkably, despite zero training on any human data, the resulting model demonstrates this human-like trend which is not observed in existing AI systems. The stereo audio used to train our model is generated using 3D simulation engines and captures the interaural distance characteristic of human ears. These findings suggest that strong audio-visual alignment between AI and human behavior stems from both the quality of sensory inputs and the physical structure of human sensory systems. We highlight our key contributions below:
1. We introduce a unified framework to systematically benchmark audio-visual localization models under modality conflicts, absence, and misalignment. We develop a scalable method to generate two-channel stereo audio for static images using 3D simulation engines, enabling the creation of diverse, naturalistic, and spatially grounded audio-visual datasets.
2. We design and conduct psychophysics experiments to assess human strategies in resolving audio-visual conflicts, providing a strong baseline for AI-human comparison. We provide a detailed analysis of modality conflicts and biases in existing state-of-the-art SSL models and humans, highlighting key performance and behavioral differences under challenging multi-modal conditions.
3. We argue that strong audio-visual alignment between AI and human behavior emerges from high-quality sensory inputs and the anatomical structure of human sensory systems. To demonstrate this, we fine-tune a state-of-the-art model on our realistic, spatially grounded, well-controlled audio-visual dataset. The resulting model not only significantly outperforms existing AI systems but also displays human-like spatial biases in localization precision.
2 Experiments
2.1 AudioCOCO dataset
Image selection. We adopt the MSCOCO image dataset coco for its wide coverage of everyday objects. To avoid overlap with data potentially seen during AI model training, we only use the 29,737 images from the test set. Since some objects are more likely to produce sound in real-world contexts, we manually selected 12 audible object categories from the MSCOCO annotations—spanning humans, animals, vehicles, and electronic devices. Recognizing that larger objects may be easier to detect and localize, we further categorize sounding objects by their relative size in the image. Object size is defined as the ratio of the object’s segmentation mask area to the total image area, independent of their real-world scale. We define three size bins: Size1 (0–5%), Size2 (5–15%), and Size3 (15–30%). Objects occupying more than 30% of the image are excluded, as they make the localization task trivially easy. To ensure class balance, we cap the number of images per object size per category at 150. For experimental control, each image contains only one target sounding object, with no other instances of the same category. For the multi-instance SSL task (see Sec 2.2), we select 150 images per object size per category, each containing 2–5 instances of the same semantic category, with exactly one designated as the sounding target. See Supp Fig. S1 and Supp Fig. S2 for the distribution of image counts by category and the spatial locations of target sounding objects.
Figure 2: Overview of congruent and manipulated vision-audio conditions in our UniAV framework and task schematic. An example of the vision-audio congruent condition (a) is shown, where a dog sound is played (black box) with a matching visual source. Visual and auditory modifications for the other five experimental conditions (b-f) are also displayed. See Sec 2.2 for more details. (h) Each trial began with a fixation cross (500 ms), followed by the presentation of an image-audio pair from either of the six conditions (a-f). Participants were instructed to use the computer mouse to click on the perceived location of the sound source within 20 seconds.
Audio selection. Prior work often pairs COCO images with external audio datasets like VGGSound vggsound or FSDnoisy18k FSDnoisy18k , but these audio sources often lack spatial and temporal consistency. For example, moving sound sources (e.g., a dog running across the video frames) cause spatial shifts that make them unsuitable for generating realistic stereo audio on static images. Additionally, recordings may include background noise or mixed sounds from multiple objects, reducing semantic clarity and overall quality. To address these issues, we apply the three filtering steps to the videos from VGGSound vggsound , ensuring that only high-quality, semantically aligned audios are retained. See Supp Sec. S1 and Supp Fig. S3 for detailed filtering steps and results. After applying these three filtering criteria, we have 720 high-quality audio clips from VGGSound vggsound . Combined with the 7,284 selected images from MSCOCO coco , we proceed to synthesize depth-aware stereo audio-image pairs.
Depth-aware stereo audio synthesis. To generate spatialized stereo audio for their corresponding sounding objects on static images, we employ Unity bomatter2021pigs as our 3D simulation engine, which allows us to control precise object locations and simulate realistic stereo sound based on the spatial layout of visual scenes. As illustrated in Fig. 1(b), we define a Cartesian coordinate system within Unity. The listener is positioned at the origin (0, 0, 0). The computer screen displays the image to listener, is placed parallel to the x–z plane and aligned along the positive y-axis, with a fixed physical distance of 0.76 meters from the listener in the human psychophysics experiment (Sec 2.3). To estimate the depth of objects within the 2D image, we utilize the DepthAnything model depthanything , which outputs a relative depth map with values ranging from 0 to 10. However, without access to the original camera intrinsics of MSCOCO images, determining absolute scene scale in Unity is nontrivial. To address this, we normalize the depth values $d_{p}$ of the image as $d_{p}^{\text{norm}}=d_{p,\max}-d_{p}$ where $d_{p,\max}$ are the maximum depth values in the image. Sound loudness decreases logarithmically with distance from the listener. To maintain sound amplitudes within a comfortable and perceptible range, we rescale the normalized depth by 0.5. The final y-coordinate in Unity $y_{u}$ for the sounding object is computed as $y_{u}=d_{p}^{\text{norm}}/2+0.76$. For the x and z coordinates of the sounding object in Unity, we map the object’s pixel location from the 2D image to the physical dimensions of the monitor. Given that the display has a resolution of 90 pixels per inch, we compute $x_{u}=x_{p}/90$ and $z_{u}=z_{p}/90$, where $x_{p}$ and $z_{p}$ are the pixel coordinates of the target object’s center in the image. Once the sounding object’s 3D position $(x_{u},y_{u},z_{u})$ is established in Unity, we place a static audio source at this location. Using an interaural distance of 0.17 meters, reflecting the typical distance between human ears, Unity simulates realistic two-channel stereo sound based on the spatial relationship between the listener and the sound source. This procedure allows us to synthesize spatially grounded, depth-aware stereo audio for each image-sound pair.
2.2 Experimental conditions
As illustrated in Fig. 2, we introduce six experimental conditions to systematically examine the modalitity biases and conflicts in humans and AI models. Each condition comprises 184,227 image-audio pairs, except for the VOnly and MultiInstLoc conditions.
Audio-visual Congruent (Congruent) represents the ideal scenario where both the audio semantics and localization align perfectly with the corresponding visual target’s semantics and location. This should serve as the upper bound for performance in both humans and AI models. For instance, as shown in Fig. 2(a), a dog sound is played at the same location as the dog in the image.
Conflicting Visual Cue (ConflictVcue) examines the scenario where the semantics of both the visual and audio cues belong to the correct category but are spatially misaligned. In Fig. 2(b), a dog sound is played at the location of a cat, while the silent dog is visually present at a different location. Among all the image-audio pairs in Congruent condition, we randomly choose an object from a non-target category as the sound source. We do not limit the distance between the distractor and the target.
Absent Visual Cue (AbsVcue) explores the case when a target sound is present but the visual scene contains non-relevant objects, and no visual cue matches the sound. For instance, in Fig. 2(c), a dog sound might be played on the cat, but no dog is visually present. From the image-audio pairs in the congruent condition, we randomly select a target sound to play at a randomly selected sounding object in the scene where no relevant objects aligning with the semantics of the sound source exists. This condition is more stringent than Conflicting Visual Cue, as it lacks any visual cues altogether.
Audio Only (AOnly) represents the extreme case where no meaningful visual information is provided, and the sound source is randomly placed anywhere within the image. The image could be a blank gray image with pixel values set to 128 (Fig. 2(d)) or a Gaussian noise image with a mean of 0 and a standard deviation of 1. For example, a dog sound could be randomly played on the left side of a pure gray background.
Vision Only (VOnly) exploits multi-modal biases or priors. In this condition, only visual scenes are provided to the AI models, and they must localize the sounding object despite the absence of any meaningful sound. The audio could either be completely silent or filled with random Gaussian noise (mean 0, standard deviation 1). For example, the same visual stimulus as in the congruent condition is presented, but the dog sound is replaced with silence or noise (Fig. 2(e)). We use the same set of 2,920 images as in the congruent condition.
Multi-Instance Localization (MultiInstLoc) follows the same motivation as the cocktail party problem haykin2005cocktail and features several objects of the same category, but the audio corresponds to only one specific instance, testing localization accuracy in a multi-instance scenario. For example, as illustrated in Fig. 2(f), a dog sound is played at the location of the left dog, while both dogs are visually present in the scene. This condition follows the same setup as the Congruent condition, but is more challenging due to the presence of multiple visually relevant objects. We selected images from the test set of the MSCOCO dataset, where the number of object instances within a given image is restricted to between 2 and 5 for the target category. This results in 263,919 image-audio pairs.
2.3 Human psychophysics experiment
We conducted in-lab psychophysics experiments with 15 participants, collecting a total of 4,050 trials. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. Every experiment lasted approximately 40 minutes. The experimental setup is schematically illustrated in Fig. 2(h). Each trial began with a fixation cross displayed for 500 milliseconds, followed by the presentation of an image and audio pair drawn from one of the six experimental conditions (see Sec 2.2). The 6-second audio clip paired with the image stimulus continuously loops until the trial concludes. Participants were instructed to use a computer mouse to click on the perceived location of the sound source within a time limit of 20 seconds, while wearing stereo headphones throughout the experiment. All trials were randomly sampled, and their presentation sequences were shuffled to minimize order effects. If participants failed to respond within the allotted 20 seconds, the trial automatically ended, and the next trial commenced. Instead of relying on the pre-rendered image-audio pairs, we conducted an audio calibration procedure at the start of the experiment to account for individual differences in height and auditory perception. To achieve this, we implemented real-time stereo audio synthesis in Unity. Following calibration, an audio validation task was conducted to ensure successful calibration and adequate spatial hearing accuracy from the participants. For more details about audio calibration and validation, see Supp Sec. S2 and Supp Fig.S4.
Figure 3: Overview of state-of-the-art multi-modal models for sound source localization (SSL). All models take a pair of images and stereo or mono audio as input, processing them separately through vision and audio encoders. For CNN-based models (indicated by blue arrows), feature maps from the final layers of each encoder are extracted, compared via cosine similarity, and used to generate a similarity heatmap. For transformer-based models, output token sequences are obtained from both encoders. Dot products on their token embeddings are calculated, and a heatmap is produced accordingly. During evaluation, SSL accuracy is determined by verifying whether the location of the maximum activation on the heatmaps (red triangles or red tokens) lies within the segmentation mask of the ground truth sounding object (red bounding box).
2.4 Methods and Evaluation Metrics
We benchmarked ten state-of-the-art multi-modal models on the same set of stimuli used in our human experiments. These models include SSLTI ssltie , LVS vggss , FNAC fnac , IS3 is3 , the fine-tuned version of IS3 (IS3-FT), CAVP cavp , AVSegformer gao2024avsegformer , ImageBind imagebind and LanguageBind languagebind . We provide a unified overview of all the SSL models in Fig. 3. We briefly introduce IS3, IS3-FT, and random in the main text. See Supp Sec. S3 for the details about other models.
IS3 is3 is a dual-stream architecture with 2D Convolutional Neural Networks (CNNs), which processes visual and auditory inputs separately using dedicated encoders before fusing the features for contrastive learning during training. IS3 also includes an Intersection-over-Union (IoU) loss and a semantic alignment loss to improve localization accuracy during supervised training. The model is trained on the FlickrSoundNet aytar2016soundnet and VGG-Sound vggsound datasets.
IS3-FT is introduced to examine how training data quality and audio input format (stereo vs. mono) affect model behavior. We fine-tune the pre-trained IS3 is3 on our AudioCOCO training set (see Fig. 3). For the mono condition (IS3-FT-Mono), we average the two stereo channels into a single-channel input, which is then passed to the pre-trained audio encoder. For the stereo condition (IS3-FT-Stereo), we add a 1D convolutional layer before the audio encoder to fuse stereo features. In both setups, all network parameters are made learnable during fine-tuning.
Random is a chance model that randomly selects a location on the image as the predicted sound source location. It serves as a lower bound for SSL without using any audio-visual information.
For both IS3 and IS3-FT models, feature maps from the final layers of the encoders are extracted, and cosine similarity is computed between them to produce a similarity heatmap. The predicted sound source location is identified as the point with the highest activation on the heatmap. See Supp Sec. S3 for how the sound source location is predicted by the transformer-based models.
Evaluation metrics. To disentangle spatial localization from semantic alignment between visual and audio modalities, we define two metrics: Audio Accuracy (A-Acc) measures whether the model or human localizes the true sound source regardless of matching semantics. A-Acc = 1 if the peak activation falls within the segmentation mask of the sounding object; 0 otherwise. Vision Accuracy (V-Acc) measures alignment with visual semantics. V-Acc = 1 if the peak activation falls within any object of the correct category, even if it might not be the actual sound source, such as in MultiInstLoc conditions. In VOnly condition, V-Acc = 1 if the activation overlaps with any object from the 12 sound-emitting categories in AudioCOCO.
3 Results
We present the results of humans and AI models across all experimental conditions. Due to space constraints, we highlight comparisons between humans and the top performing AI models IS3 and IS3-FT-Stereo. Results for the remaining models are provided in Supp Tab. S1.
Figure 4: Object size matters for humans and AI models in the congruent condition. Accuracy increases with object sizes for both humans and AI models, with humans outperforming AI models, especially for small targets. Here and in subsequent figures, error bars represent Standard Error Mean (SEM).
Figure 5: ConflictVCue and AbsVCue hurt SSL performance for both humans and AI models. The accuracy of humans and AI models under ConflicVcue and AbsVcue conditions is shown for object size 2, with results from the congruent condition included for comparison. See Supp Tab. S1 for the results on object sizes 1 and 3.
Object size matters for humans and AI. As shown in Fig. 5, human A-Acc improves steadily with increasing object size, indicating that larger sounding objects are easier for humans to localize. Similarly, IS3-FT-Stereo shows a notable improvement in A-Acc from 19.3% to 52.2% as object size increases, though it still lags behind human performance. Among all models, IS3-FT-Stereo consistently outperforms the second-best model, IS3, across all object sizes. This is likely because IS3 is trained on datasets biased toward large, centered objects, whereas AudioCOCO offers greater diversity in sizes of sounding objects and their spatial locations. Despite these differences, both humans and all AI models achieve above-chance A-Acc when localizing large objects. However, only humans and IS3-FT-Stereo maintain consistently strong performance across all object sizes, including small objects, whereas IS3 does not.
Figure 6: Humans and stereo-trained models show horizontal-vertical asymmetry in SSL accuracy, while mono-trained models do not. We report the proportion of trials (object size 2) where the predicted sound location falls within 6 degrees of visual angle from the ground truth, separately along the horizontal (textured bars) and vertical (plain bars) directions. See Supp Tab. S1 for results on object sizes 1 and 3.
Figure 7: AI models show a bias towards objects that emit sounds, even in the absence of sound or in the presence of noisy sound. The V-Acc of AI models is presented in the VOnly condition, where either no sound or only noisy sound is present. Despite the lack of auditory cues, V-Acc of AI models on object size 2 remains higher than random (dark gray), indicating a bias towards audible objects.
Conflict cues harm more than the lack of cues for AI. As shown in Fig.5, ConflictVCue and AbsVCue impair SSL performance for both humans and AI models, compared to the congruent conditions. While humans and IS3-FT-Stereo can still outperform chance, IS3 drops to near chance performance under these conditions. This suggests that humans and IS3-FT-Stereo are more robust to SSL with conflicting or missing visual cues. Interestingly, for IS3-FT-Stereo, conflicting visual cues cause a greater performance decline than the absence of visual cues that align with the sound source. This implies that IS3-FT-Stereo can be misled by conflicting visual cues. These findings suggest that, unlike humans, current AI models still place greater reliance on visual input for SSL.
AI fails when there are only auditory cues but humans can. In Supp Fig. S5, humans achieve A-Acc above chance even without any visual information (e.g., grayscale or Gaussian noise backgrounds), confirming that humans can perform SSL based on auditory cues alone. However, their performance is notably lower than in the congruent condition, highlighting that congruent visual information enhances SSL for humans. In contrast, IS3-FT-Stereo and IS3 performs worse than humans and drops to chance-level performance in AOnly condition, recapitulating that they depend heavily on congruent visual cue to support SSL.
In neuroscience francl2022deep , horizontal-vertical asymmetry in auditory precision describes the tendency for better localization in the horizontal than vertical axes. We measured the proportion of trials based on the Euclidean distance between predicted and ground-truth sound locations along each axis. As shown in Fig. 6, humans show a strong horizontal bias, with 52.1% of predictions falling within 6 degrees of visual angle horizontally, compared to 31.5% vertically. Remarkably, IS3-FT-Stereo—despite no exposure to human data—exhibits a similar, though weaker, bias. In contrast, IS3 shows no clear directional preference, likely reflecting its limited use of spatial stereo cues.
AI biases towards sound-emitting objects. Previous studies cavp ; is3 report that AI models often rely on visual shortcuts—favoring large, centered objects—due to dataset biases. To probe additional behavioral tendencies, we evaluate models under a vision-only condition. Without meaningful sound, any object could plausibly be the source. As shown in Fig. 7, models tend to favor sound-emitting objects (e.g., people, animals) over irrelevant ones (e.g., sky, trees), yielding significantly above-chance V-Acc. This suggests models carry prior knowledge about which objects typically produce sound. Interestingly, even noise input increases V-Acc compared to silence, indicating that audio noise may enhance this bias for AI. IS3-FT-Stereo exhibits the strongest effect, with substantially higher V-Acc than IS3.
(a) Multi-Instance Localization    (b) Overall Performance
Acc(%) Rand
IS3
is3
CAVP
cavp
AVSeg
gao2024avsegformer
SSLTI
ssltie
FNAC
fnac
IS3-FT
-Stereo
Human    Acc(%) Rand
IS3
is3
IS3-FT
A- Acc Size1 1.6 4.2 5.9 5.5 0.6 0.7 8.9 29.1  Mono Size1 1.6 9.8 13.1
Size2 9.1 10.3 11.5 11.3 2.5 3.9 14.3 63.9  Size3 9.5 17.9 25.6
Size3 21.3 22.4 27.4 26.5 24.0 26.3 31.5 71.1  Size3 21.8 35.0 44.3
V- Acc Size1 8.4 17.2 20.5 19.6 13.2 13.4 85.4 97.5  Stereo Size1 1.6 11.5 17.5
Size2 17.8 22.8 26.0 24.7 20.7 25.5 89.5 99.1  Size2 9.5 21.3 30.0
Size3 26.2 49.5 59.5 56.3 49.4 52.8 93.7 99.8  Size3 21.9 42.0 45.7
Table 1: Multi-instance SSL remains a challenging task for both humans and AI models. The table on the left summarizes the audio and visual localization accuracy of humans and AI models under the multi-instance condition across all object sizes. For AI models, both input data quality and the use of stereo audio substantially impact SSL performance. The table on the right reports the overall SSL performance in Audio Accuracy (A-Acc) of AI models trained with either mono or stereo audio described in Sec.2.2, excluding the VOnly condition. In both tables, bold and underlined values indicate the best and second-best performances, respectively.
Multi-instance SSL remains challenging for humans and AI. As shown in Tab. 1a, V-Acc is high for both IS3-FT-Stereo and humans, indicating strong alignment between audio and visual semantics. This allows them to use audio cues to identify all visual objects with matching semantics. However, in scenes with multiple object instances, successful localization requires fine-grained SSL, beyond semantic matching. In these cases, A-Acc drops for both humans and IS3-FT-Stereo. Compared to the Congruent condition (with a single sounding object), multi-instance SSL is notably harder, as it demands precise spatial disambiguation. Despite this, both IS3-FT-Stereo and humans still perform above chance, especially for small objects. This demonstrates an ability to localize sounds at fine spatial resolution for both humans and IS3-FT-Stereo. Nonetheless, IS3-FT-Stereo still lags behind human performance, with the gap more pronounced in A-Acc than in V-Acc, underscoring the need for improved SSL in AI. Among all models, IS3-FT-Stereo performs best. It even outperforms strong baselines like CAVP and AVSeg, which are trained on large-scale, standard audio-visual datasets.
Training data quality and stereo input are important for SSL. We report the average A-Acc across all object sizes for Random, IS3, and two fine-tuned variants of IS3 over five experimental conditions in Tab. 1b and for all other AI models in Supp Tab. S1 and Supp Sec. S4. IS3-FT-Stereo consistently outperforms all baselines by a large margin. Tab. 1 also shows that IS3-FT-Mono outperforms the original IS3 even with fewer examples to fine-tune on, highlighting the impact of training data quality. Moreover, IS3-FT-Stereo surpasses IS3-FT-Mono, underscoring the importance of stereo inputs in capturing spatial localization cues.
We present qualitative results in Fig. 8, comparing predicted sound source locations from humans, IS3, and IS3-FT-Stereo. IS3-FT-Stereo localizes sounds more accurately and often matches human predictions. In contrast, IS3 struggles with small or off-center objects. For example, IS3 fails to localize a dog in the top-right of an image in the Congruent condition and it frequently mislocalizes sources in incongruent conditions, such as horses, dogs, or cats. This suggests IS3 overly relies on visual cues and fails when auditory signals dominate. Nonetheless, IS3-FT-Stereo still underperforms humans in some cases. Additional qualitative examples are shown in Supp Fig. S6, indicating that IS3-FT-Stereo lacks the flexibility to resolve audio-visual conflicts or rely solely on audio when visual cues are absent.
Figure 8: IS3 struggles to localize sound sources, whereas humans and IS3-FT-Stereo perform well across all four experimental conditions. The leftmost images (Columns 1 and 4) show the correct localization results made by human participants. Red boxes mark the ground truth sound source locations, while green circles indicate mouse click responses from multiple participants. The middle columns (2 and 5) and rightmost columns (3 and 6) display heatmaps predicted by IS3-FT-Stereo and IS3, respectively. Yellow triangles on the heatmaps denote the predicted sound source locations from each model. See the colorbar for the activation values of the heatmaps.
4 Discussion
We systematically and quantitatively examine modality biases and conflict resolution in sound source localization (SSL) across humans and AI models. Our study spans six audiovisual conditions, including congruent cues, conflicting audio-visual signals, and scenarios with missing audio or visual input in natural scenes. Our results reveal that object size significantly affects human performance. While conflicting or absent visual cues impair human SSL, people remain robust at successfully localizing even small sounding objects under challenging conditions, such as in multi-instance settings or with no visual input at all. In contrast, current multimodal AI models show strong dependence on visual cues. They often misattribute sound to large, centered, and audible objects and suffer sharp performance drops when visual cues are absent. Conflicting visual cues can even further degrade the model performance. Most models perform near chance when localizing small sound sources without accompanying visuals.
We hypothesize that these limitations stem from two key issues: the use of low-quality, biased audiovisual datasets and monaural audio inputs that lack spatial fidelity. To address this, we curate a high-quality dataset—AudioCOCO—by applying rigorous filtering criteria to audio and images. Using depth maps and 3D simulation engines, we generate realistic, spatialized stereo audio that obeys the physical laws of sound propagation. We fine-tune existing models on AudioCOCO using stereo inputs. Despite using fewer training examples, the resulting model surpasses current state-of-the-art across all conditions. Remarkably, without training on human behavior, the resulting model displays human-like localization biases, favoring horizontal over vertical precision. Ablation studies confirm the critical role of both dataset quality and stereo input in capturing spatial cues. This work highlights the importance of curating less biased datasets and building models that respect the constraints of physical sensory systems. Future directions include scaling up the dataset, incorporating temporal dynamics from video, and enhancing realism in simulated sound rendering.
Appendix S1 Details about the AudioCOCO Dataset
(a) The spatial distribution of single-object by category.
(b) The spatial distribution of multi-objects by category.
Figure S1: The visualization results of images’ spatial distributions in the AudioCOCO dataset are presented for each object category, where the color bar indicates the frequency of object occurrences at each spatial location within the images. Darker regions correspond to higher frequencies.
(a) The count distribution of multi-objects by category.
(b) The count distribution of single-object by category.
Figure S2: The visualization results of the image count distributions in the AudioCOCO dataset are presented, where blue, green, and red represent object size1, size2, and size3, respectively. This visualization illustrates how instances are distributed across different object size categories, highlighting the dataset’s balanced distribution in terms of object sizes.
More details on image selection. The COCO 2014 dataset offers annotations for 80 object categories, including both bounding boxes and segmentation masks. To ensure a fair comparison—especially given that some model backbones are pre-trained on COCO—we utilized the COCO 2014 validation set rather than the training set. From this, we selected 12 common vocal categories frequently encountered in daily life: person, motorbike, train, boat, elephant, bird, cat, dog, horse, sheep, cow, and keyboard. A total of 29,737 images containing at least one vocal object were extracted to form the pool of image candidates for our dataset.
To investigate spatial distribution, we visualized the occurrence frequency of these categories across the 29,737 images in Fig. 1(b) and Fig. 1(a), grouping instances by three object area size bins defined in the main draft: object size1 (0-5%), size2 (5-15%), and size3 (15-30%). For smaller object sizes, category locations were relatively uniformly distributed across the image space. However, for larger objects, some categories—such as bus, train, and truck—exhibited a strong center bias. To correct for this, we filtered the dataset to ensure that no position in the final heatmap exceeded 50% in frequency.
Additionally, given that the image distribution over all object categories in COCO is not uniform, we performed random sampling to cap the number of instances to 150 per category per object size bin, promoting balanced representation within the AudioCOCO dataset. The final distributions after filtering are shown in Fig. 2(a) and Fig. 2(b), demonstrating that AudioCOCO achieves a balanced dataset across object area sizes, object center positions, and categories.
More details on audio selection. To ensure that only high-quality, semantically aligned clips are retained, we apply the following three filtering steps to the videos from VGGSound vggsound . First, we introduce Semantic Consistency (SeC) and select audios that are representative of the semantic object categories. Specifically, we take all audio files belonging to the same semantic category, extract audio features from the last layer of the Wav2vec baevski2020wav2vec model, compute their pari-wise cosine similarities based on these features, and retain the top 80% audios with the highest cosine similarity scores. Second, to ensure the audio is free from noise or interference from unrelated sources, we introduce Mel-Spectrogram Similarity (MSS) as a filtering criterion within each category. For each audio clip, we compute its Mel spectrogram, average the frequency magnitudes over time, and apply a logarithmic transformation to compress high-frequency components—yielding a compact representation of the audio’s overall spectral structure. We then calculate the cosine similarity between these representations and retain the top 65% of audio clips based on this MSS metric, following the SeC criteria. Third, to ensure the stereo audio corresponds to the sounding object being centered in the video frame, we introduce Spatial Consistency (SpC) for each audio-image pair. We compute the Spearman correlation between the left and right audio channels; a high correlation suggests the sound source is centrally located, producing similar waveforms in both channels. We retain the top 50% of audio clips based on this metric from the MSS-filtered set.
For audio selection, we also enforce a minimum threshold of 20 audio clips per category. If any filtering step results in a category falling below this threshold, that specific step is discarded, and the previous filtering output is retained as the final result to ensure sufficient data coverage across all categories. The final distributions after filtering are shown in Fig. S3.
(a) The three stages of audio filtering result.
(b) The comparison between original audio counts and filtering audio counts
Figure S3: The visualization results of audio statistics in the AudioCOCO dataset demonstrate that the audio quality across most categories has significantly improved following the filtering process. Additionally, we ensured a balanced count distribution among all categories. These outcomes highlight the effectiveness of our selection and refinement strategy in enhancing both the semantic consistency and acoustic quality of the dataset.
Appendix S2 Details of human psychophysics experiments
All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. Participants were instructed to use a computer mouse to click on the perceived location of the sound source within a time limit of 20 seconds, while wearing stereo headphones (SENNHEISER MOMENTUM 4 with active noise cancellation) throughout the experiment.
Audio Calibration. Instead of relying on the pre-rendered image-audio pairs, we conducted an audio calibration procedure at the start of the experiment to account for individual differences in height and auditory perception. To achieve this, we implemented real-time stereo audio synthesis in Unity, which communicates with MATLAB (hosting the human behavioral experiment) via TCP connections. The measured delay for real-time stereo sound synthesis and presentation is within 500 milliseconds, ensuring seamless interaction between the two systems.
During calibration, participants were presented with a white dot at a random location on the screen, accompanied by an audio clip spatially rendered at the dot’s position. A white cross at the center of the screen served as a spatial reference (see Supp Fig.S4). Using the keyboard’s arrow keys, participants adjusted the perceived audio source position until the sound aligned with the white dot. Once satisfied, participants pressed the ESC key to confirm the alignment. This process was repeated six times with different white dot positions.
For each calibration step $t\in{1,2,\dots,6}$, we recorded the participants’ adjustments of the sound source location in pixels $(\Delta x_{t},\Delta z_{t})$. The calibration hyperparameters $\alpha$ and $\beta$ were computed as the mean of $\Delta x_{t}$ and the sum of $\Delta z_{t}$, respectively. These hyperparameters were then applied to scale the Unity coordinates $x_{u}$ and $z_{u}$, correcting for individual perceptual biases in auditory localization during the main experiment. The rationale for using the mean adjustment for $\Delta x_{t}$ (horizontal direction) and the sum for $\Delta z_{t}$ (vertical direction) is based on human spatial hearing characteristics—listeners typically localize horizontal (azimuth) sound sources with higher precision than vertical ones. This design allows greater tolerance for variability in the vertical dimension (altitude) during calibration, ensuring more robust alignment with participants’ perceptual expectations.
Audio Validation. Following calibration, an audio validation task was conducted to ensure successful calibration and adequate spatial hearing accuracy. Participants heard an audio clip played from one of two possible locations on a horizontally arranged 1x2 grid displayed on the screen (see Supp Fig. S4). They were instructed to click on the half of the screen from which they perceived the sound. This validation procedure was repeated six times. If a participant’s spatial sound localization accuracy fell below 83%, the calibration and validation procedures were repeated to guarantee reliable data quality during the actual experiment.
Figure S4: The audio calibration and validation process. During calibration, the player is at a random position close to the white dot. The red arrow represents a potential movement trajectory. During validation, the display is divided into two equal quadrants, with the audio player appearing at a random position three times within each quadrant.
Appendix S3 More Implementation Details of AI Models
SSLTI ssltie , LVS vggss , FNAC fnac , and IS3 is3 are SSL models based on dual-stream architectures with 2D Convolutional Neural Networks (CNNs). Each model processes visual and auditory inputs separately using dedicated encoders before fusing the features for contrastive learning during training. Both encoders are based on ResNet18 resnet . Beyond standard contrastive learning, IS3 introduces an Intersection-over-Union (IoU) loss and a semantic alignment loss to improve localization accuracy and better alignment between audio and visual modalities during supervised training. All the models are trained on the FlickrSoundNet aytar2016soundnet and VGG-Sound vggsound datasets.
CAVP cavp is a sound source segmentation model that follows a dual-stream CNN architecture. It uses PVTV2-B5 wang2022pvt as the visual encoder and either VGGish vggish or ResNet18 as the audio encoder. CAVP is trained in a fully supervised manner using cross-entropy loss for segmentation and contrastive learning to align audio-visual features. Training data includes AVSBench zhou2024audio and VPO cavp datasets.
AVSegformer gao2024avsegformer is an audio-visual semantic segmentation model built on a dual-stream transformer-based architecture. It employs SegFormer xie2021segformer as the visual backbone and Audio Spectrogram Transformer (AST) ast as the audio encoder to capture fine-grained cross-modal representations. AVSegformer integrates both modality-specific and fused token embeddings through a lightweight fusion decoder for pixel-level prediction. It is trained in a fully supervised setting using a combination of cross-entropy loss for segmentation and audio-visual consistency loss. AVSegformer’s training data includes three subsets of AVSBench.
ImageBind imagebind and LanguageBind languagebind are large-scale, multi-modal transformer models with billions of parameters. These models are trained with contrastive learning objectives across six modalities: image, audio, video, text, depth, thermal, and IMU, and embed these into a shared representation space. Both models use a Vision Transformer(ViT) vit for visual encoding and AST ast for audio encoding. ImageBind is trained on large-scale datasets including LAION laion ; cherti2023reproducible , SSv2 ssv2 , and K400 k400 . LanguageBind builds on the pre-trained encoders from ImageBind and further fine-tunes them on the VIDAL-10M languagebind dataset.
Figure S5: While humans benefit from visual cues, they remain capable without them in the AOnly condition, unlike AI models, which fail to localize sound sources without congruent visual information. The accuracy under AOnly conditions with both grayscale and Gaussian noise backgrounds drops for both humans and AI models compared to the congruent condition on object size 2.
Figure S6: The figure shows the failure examples of both IS3 and IS3-FT-Stereo compared to humans. The leftmost images (Columns 1 and 4) show the correct localization results made by humans. Red boxes mark the ground truth sound source locations, while green circles indicate mouse click responses from multiple participants. The middle columns (2 and 5) and rightmost columns (3 and 6) display heatmaps predicted by IS3-FT-Stereo and IS3, respectively. Yellow triangles on the heatmaps denote the predicted sound source locations from each model. See the colorbar for the activation values of the heatmaps.
Implementation details of AI models. All ten models—except IS3-FT-Stereo—take paired image and mono audio inputs, processing each modality independently through dedicated visual and audio encoders. For the mono audio setup, we follow the IS3-FT-Mono design by averaging the stereo channels into a single-channel input. For CNN-based models, feature maps from the final layers of the encoders are extracted, and cosine similarity is computed between them to produce a similarity heatmap. For transformer-based models, token sequences from both encoders are retrieved, and pairwise dot products of their token embeddings are calculated to generate the heatmap. The predicted sound source location is identified as the point with the highest activation on the heatmap.
We evaluated all the current models using their publicly available, pre-trained weights and adhered to their original implementation details. Since AVSBench zhou2024audio shares some images with our AudioCOCO test set, we exclude those overlapping image-audio pairs when evaluating CAVP cavp and AVSegformer gao2024avsegformer . Experiments for IS3-FT, ImageBind and LanguageBind were conducted on 8 NVIDIA A100 GPUs, while all other models were tested on 4 NVIDIA A6000 GPUs. To fine-tune IS3-FT models, we use the SGD optimizer with a momentum of 0.9 and weight decay of 0.0005. The initial learning rate is set to 0.001, and the batch size is 32. All the model evaluations are run 3 times with different random seeds.
Appendix S4 More Qualitative Results of Humans and AI in SSL
Congruent Conflicting VCue Absent VCue Audio Only Vision Only Multi-Instance Model Acc(%) A-Acc Vision Gray Vision Noise Audio Silent Audio Noise A-Acc V-Acc Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Random 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.1 21.3 8.4 17.8 26.2 IS3is3 4.3 13.5 31.0 4.3 7.9 14.5 6.2 11.0 20.8 1.4 5.2 9.4 1.5 5.4 10.7 25.9 32.0 69.5 30.3 58.0 77.6 4.2 10.3 22.4 17.2 22.8 49.5 FNACfnac 3.8 11.5 26.9 3.9 7.2 13.3 5.8 10.5 18.6 1.1 4.8 8.7 1.3 5.0 9.6 23.2 26.5 63.8 27.4 52.6 71.9 0.7 3.9 26.3 13.4 25.5 52.8 SSLTIEssltie 3.4 9.9 23.7 3.1 6.8 12.4 5.2 9.9 16.5 0.7 3.9 8.2 1.1 4.3 8.9 23.1 26.5 60.7 25.5 49.7 69.8 0.6 2.5 24.0 13.2 20.7 49.4 LVSvggss 3.3 9.7 23.6 3.0 6.5 12.1 5.0 9.4 16.0 0.7 3.8 8.3 1.0 4.2 8.8 23.2 26.4 60.5 25.4 49.7 70.2 0.5 3.9 25.3 12.5 22.4 50.6 CAVPcavp 5.1 15.2 38.7 3.8 7.0 13.2 5.5 10.2 18.4 0.9 4.7 8.4 1.2 4.9 9.8 25.4 30.7 70.5 29.4 56.5 72.3 5.9 11.5 27.4 20.5 26.0 59.5 AVSegformergao2024avsegformer 4.7 15.5 39.5 3.6 7.1 12.8 5.3 9.8 17.2 0.9 4.5 7.9 1.1 4.6 9.4 25.0 30.3 69.9 28.7 55.8 70.6 5.5 11.3 26.5 19.6 24.7 56.3 ImageBindimagebind 1.4 5.8 14.5 2.2 5.7 11.3 3.4 7.2 12.8 0.3 3.1 6.6 0.4 3.8 7.5 19.8 24.5 47.2 21.2 26.3 52.0 0.5 2.7 15.3 6.5 12.4 25.1 LanguageBindlanguagebind 1.1 4.6 12.7 2.1 5.4 10.7 3.2 6.9 10.6 0.2 3.0 6.5 0.4 3.7 7.6 18.6 23.2 44.9 19.1 23.8 48.8 0.4 2.1 13.0 6.2 10.8 22.5   IS3-FT -Stereo 19.3 29.1 52.2 8.8 14.7 27.5 10.2 18.2 31.0 3.1 12.3 23.7 3.8 13.5 26.0 40.6 61.1 83.2 45.5 76.8 90.5 8.9 14.3 31.5 85.4 89.5 93.7 Human 32.9 56.6 69.3 22.8 39.5 47.1 12.7 35.3 41.2 4.5 30.3 45.6 4.2 28.8 44.9 - - 18.2 30.5 52.8 97.5 99.1 99.8 Table S1: Comparison of AI models in UniAV. Bold and underlined values indicate the best and second-best performances, respectively.
As illustrated in Fig. S6, we further highlight the limitations of IS3-FT-Stereo in comparison to human performance. Specifically, IS3-FT-Stereo struggles to localize the correct sounding object when a distractor occupies a large visual region in the scene. This suggests that the model is heavily influenced by visual saliency, often prioritizing visually prominent objects over auditory cues, a bias that humans are better able to overcome.